# Establishing Reality-Virtuality Interconnections in Urban Digital Twins for Superior Intelligent Road Inspection

### v1

@@ IV-C Digital Road Twin Generator
with similar solutions for and . With all intersecting faces removed, the defect models are positioned within the holes of the planar road mesh. The structure of the defect models remains fixed, except at the boundaries, which are assigned pseudo-height values. The remaining planar road surfaces must then be re-organized, typically through triangulation, which iteratively divides edges, samples vertices, and generates faces. However, as previously mentioned, mismatches may occur at the boundaries if adjacent meshes are reconstructed separately. To address this issue, the previously extracted boundaries are used as edge constraints to reconstruct the 2D structure of the road surface mesh using the constrained Delaunay triangulation algorithm [52] Since the pseudo-height is set to zero, the defect model can be seamlessly aligned with the planar road surface, ensuring geometric consistency. For better realism and diversity, road surface elevations are restored from our model library. The height of each vertex is efficiently queried through a KD-Tree built from the reconstructed surface model. The final integrated 3D mesh is then split into individual assets, as required by the simulator, to provide semantic references.
@@
V Experimental Results
V-A Simulation Scene Setup
Based on sensor data collected from defected roads, our hierarchical road model creator generates a road model library containing 53 distinct road defect models with varying shapes and sizes. These models are integrated into city scenes within the simulator, replacing the original planar road surfaces. Specifically, 646 road segments are extracted from Town01. Subsequently, 1,000 defect instances are sampled and integrated into the scene using our digital road twin generator, with randomized scales, locations, and orientations. To accelerate the generation process, we leverage parallel processing. Using 24 13th Gen Intel(R) Core(TM) i7-13700K processors and 128 GB of RAM, the integration of road meshes for Town01 is completed in approximately 2 minutes. Finally, in Unreal Engine-based simulators, semantic references are accessed via asset paths. Therefore, the generated meshes are disassembled into individual entities: the 646 road meshes replace those in the original “road” group in Carla, while the 1,000 road defects are imported into a newly created “defect” group.
V-B Perception Tasks
We design experiments for various perception tasks to demonstrate how synthetic data alleviates the scarcity of real-world road defect data and enhances of existing models.
V-B1 Semantic Segmentation
We first evaluate single-modal semantic segmentation networks using real-world RGB images. The UDTIRI dataset [7], a real-world road defect dataset containing 1,000 semantically annotated images, is used for evaluation. Each image is captured from a unique perspective using various cameras. To replicate this diversity in the simulator, virtual sensor poses are randomly generated for each capture. Experiments are conducted on a single NVIDIA RTX 4090 GPU, evaluating six single-modal networks for image segmentation and road defect detection. All networks are pre-trained for 40,000 epochs using different ratios of synthetic data, followed by fine-tuning on 600 images from the UDTIRI dataset. Input images are cropped to pixels. Although the dataset provides multiple object classes, only “background” and “pothole” are considered for this task. Network configurations follow the default settings of the MMSegmentation benchmark [53]. For example, Mask2Former [54] is trained using the AdamW optimizer [55] with a polynomial learning rate decay strategy [56]. The learning rate starts at with a weight decay of . Mean intersection over union (mIoU) is used as the evaluation metric to assess network performance.
@@ V-B1 Semantic Segmentation
| Mask2Former[54] | 77.08 | 79.79 | 80.07 |
| PSPNet[57] | 71.42 | 71.44 | 74.29 |
| SegFormer[58] | 75.16 | 75.96 | 76.94 |
| SegNext[59] | 79.68 | 79.84 | 80.63 |
| Twins[60] | 78.40 | 79.10 | 79.06 |
| UperNet[61] | 70.97 | 72.54 | 73.89 |
@@
| RTFNet [62] | 67.5 | 74.6 | 69.2 | 63.6 |
| FuseNet [63] | 66.6 | 78.7 | 53.9 | 58.5 |
| MFNet [64] | 72.0 | 80.1 | 90.6 | 72.3 |
| SNE-RoadSeg [65] | 67.4 | 78.8 | 94.5 | 73.5 |
| RoadFormer [66] | 90.3 | 94.8 | 96.1 | 90.6 |
Qualitative and quantitative results are presented in Fig. 6 and Table I, respectively. It is evident that pre-training with synthetic data significantly enhances segmentation accuracy, providing more precise boundaries and reducing the occurrence of false-positive regions. The results show that synthetic data generated by our UDT system positively impacts image-based road defect detection, increasing mIoU by up to 3%. Although the performance of Twins pre-trained with the full set of synthetic data slightly decreases 0.04 compared to the 50% subset, it still outperforms the non-pretrained model by 0.66%. This improvement alleviates the challenge posed by the limited availability of annotated real-world road defect data for road inspection models.
Single-modal prediction results from RGB images are often affected by environmental factors, such as lighting conditions. In contrast, multi-modal fusion is widely recognized as a method to enhance perception robustness, though capturing these modalities in the real world remains challenging. By leveraging the G-buffer in the rendering pipeline, we evaluate auxiliary modalities for road inspection, focusing on image segmentation for drivable areas, road defects, and backgrounds. An autopiloted vehicle is deployed in a simulated road defect scene with randomized surrounding vehicles and pedestrians. Sensors are mounted on the ego vehicle and oriented at a pitch angle of 30 degrees toward nearby road surfaces, capturing normal, depth, and event images. During data recording, dynamic weather conditions are simulated, continuously changing to enhance data variability. Since no real-world dataset provides such comprehensive information, the models are evaluated using synthetic data.
| Method | PEP(%) | EPE(pixel) | SSIM | MSE | PSNR(dB) | |
| PSMNet [67] | 53.50 | 22.00 | 2.57 | 0.82 | 152.2 | 28.3 |
| PSMNet+synthetic | 7.61 | 3.59 | 0.67 | 0.90 | 93.5 | 30.9 |
| AANet [68] | 37.60 | 23.7 | 4.79 | 0.85 | 118.2 | 29.7 |
| AANet+synthetic | 9.36 | 5.30 | 0.61 | 0.91 | 93.4 | 30.8 |
| LacGwc [69] | 15.00 | 8.17 | 1.35 | 0.89 | 113.4 | 30.4 |
| LacGwc+synthetic | 6.11 | 2.92 | 0.55 | 0.90 | 96.0 | 30.7 |
| IGEV [70] | 8.14 | 4.99 | 0.83 | 0.90 | 102.9 | 30.8 |
| IGEV+synthetic | 3.17 | 1.59 | 0.23 | 0.91 | 92.5 | 30.9 |
| ViTAStereo [71] | 7.61 | 5.08 | 1.31 | 0.90 | 124.4 | 29.6 |
| ViTAStereo+synthetic | 3.14 | 1.61 | 0.22 | 0.91 | 91.6 | 30.9 |
@@
As shown in Table II and Fig. 7, each additional modality is fused with RGB images in an RGB+X format. RGB+Depth provides the most generalized results, outperforming those trained using only RGB images by 4.5% to 12.1%. It also generates the most accurate true-positive predictions, demonstrating robust performance across all cases. This result is intuitive, as depth images directly capture the surface structure. Normal maps, which are highly sensitive to object surfaces, improve performance by 1.7% to 27.1% across most networks, except for FuseNet. Although useful in many cases, this sensitivity also introduces noise, leading to convergence issues. Fig. 7 shows that with the normal modality, FuseNet predicts the sidewalk as a noisy drivable area, resulting in poor mIoU performance. Simple fusion of event images across network channels is unstable. Although event images improve SNE-RoadSeg by 6.1%, they degrade FuseNet’s performance by 8.1%. This instability is likely due to event data being influenced by the vehicle’s driving speed, whereas RGB images assume static objects during exposure. To effectively utilize this modality, additional speed priors are required.
V-B2 Stereo Matching
For stereo matching evaluation, we compare five SoTA networks to demonstrate the improvements enabled by our system. Similar to the single-modal segmentation experiments, these networks are pre-trained on synthetic data and fine-tuned using real-world data. Unlike individual road surface images, which can still be obtained through crowdsourcing applications, stereo images of road surface defects are exceedingly rare. To address this limitation, we fine-tune the networks using the KITTI dataset [72], a widely used driving dataset that lacks road defects.
Several metrics are selected for evaluation, including percentage of error pixels (PEP), which represents the percentage of incorrect disparities with respect to a tolerance of pixels, and end-point error (EPE), which measures the average disparity estimation error. Considering the domain gap, networks fine-tuned on KITTI may struggle with road defects. To further assess accuracy, mean squared error (MSE), structural similarity index (SSIM), and peak signal-to-noise ratio (PSNR) are compared between the predicted and ground truth disparity maps. The loss function, learning rate, and optimizer settings are consistent with those reported in the corresponding publications [67, 68, 69, 70, 71].
Tables III presents the quantitative results tested on synthetic data. Nearly all networks pre-trained with our synthetic data demonstrate performance improvements. EPE is improved across all networks, with reductions ranging from 0.60 to 3.18. Although the improvement in PSNR is relatively modest, ranging from 0.1 to 2.6, qualitative results in Fig. 8 clearly show the noise difference. Networks trained solely on KITTI data perform poorly, often failing in challenging cases, such as the black areas predicted in the third image by PSMNet and AANet. In contrast, pre-trained networks produce smoother and more accurate predictions. These results confirm that our UDT system significantly enhances model performance for stereo matching tasks.
V-C Decision Making Tasks
@@ V-C1 Path Planning
In scenarios with ample road space and minimal opposing traffic, avoiding defects by following an alternative path is generally the preferred strategy. Unlike traditional obstacle constraints, road defects primarily impact the wheels rather than the vehicle body, allowing the vehicle to either bypass or glide over the defects. Consequently, collision detection at each time step is replaced by wheel-based checks. As shown in Fig. 9 (a), multiple defects are distributed along the road. The drivable area and defected regions are modeled as grid maps with a resolution of 0.2 m, and the vehicle model assumes a width of 2.0 m and a rear-to-front wheel distance of 3.0 m. Four widely used planning algorithms are compared: A* [73], RRT* [74], Lattice [75], and Hybrid A* [76]. The RRT* algorithm uses an expansion distance of 1.0 m. For the Lattice algorithm, lateral and longitudinal step sizes are set to 1.0 m and 5.1 m, respectively, with each node connecting to three adjacent nodes via a cubic spline curve. In Hybrid A*, a yaw grid resolution of 15 degrees is applied, and penalties for steering and lane changes are incorporated into the heuristic cost. Evaluation metrics are adapted from classical obstacle avoidance tasks. Path deviation quantifies the length difference between the planned path and the shortest possible path, typically the road’s midline. Path smoothness is calculated by averaging the angle changes along the trajectory, while obstacle clearance measures the mean distance between the vehicle’s wheels and the nearest road defects.
Quantitative and qualitative results of the generated paths are presented in Table IV and Fig. 9 (a). Notably, A* and RRT* do not account for vehicle kinodynamic constraints, resulting in relatively poor path smoothness, with values of 0.294 and 0.402, respectively. However, their discrete path points remain topologically reasonable. The paths generated by Hybrid A* and Lattice glide over defects and , which is justified in the given context. As one of the most commonly used planning algorithms in road scenarios, Lattice generates the shortest and smoothest path, with a path deviation of 0.87 and a smoothness value of 0.047. However, because its grids and curves are pre-sampled, obstacles may not always be perfectly avoided. In contrast, Hybrid A* expands nodes iteratively at each step, achieving an obstacle clearance of 3.74, which ensures better safety distances. The trade-off, however, is that heuristic terms must be manually designed to suit specific scenarios.
@@
V-C2 Speed Control
In scenarios where changing lanes is risky, the vehicle must traverse road defects directly. In such cases, longitudinal speed becomes the only controllable variable to minimize vibrations. We customize these scenarios using our UDT system, where surface models of real-world roads are cropped into 40-meter sections, and integrated with randomly distributed road defects. With the aid of the simulator, road defects along the wheel trajectories can be precisely located, enabling informed speed control strategies. To measure vehicle vibrations caused by defected road surfaces, the collision settings for road entities in the simulator are configured to ”use complex as simple,” allowing collisions to be calculated for each triangular face between the road and tires during physical simulation. As shown in Fig. 9 (b), six-axis IMU data is recorded as the vehicle traverses defected road sections at a constant speed. During the process of a wheel falling into and exiting a road defect, it can be observed that the most affected parameters are vertical acceleration , rotational velocity in roll and pitch . Subsequently, the vibration degree is defined as , where is set to 0.1 in this scenario. Quantitative results for traversing three road defects at different speeds are presented in Fig. 9 (c). The results indicate that, for large and severe defects such as , slowing down is the most effective strategy for driving comfort. Interestingly, for small and shallow defects, such as , increasing speed can also reduce vibrations. Although initially counterintuitive, this observation aligns with practical experience upon closer analysis.
VI Conclusion and future work
In this paper, we introduce a UDT system that leverages reality-virtual interconnections of road entities to advance intelligent road inspection. The system begins with the design of sensor equipment to capture data from defected road surfaces. A hierarchical road model creator is then used to generate detailed models of road defects and surfaces. These models are subsequently integrated into simulation scenes by a digital road twin generator, replacing the simplified planar roads commonly used in existing simulators. The proposed system provides unlimited, multi-modal, and well-annotated data for training road inspection networks, as well as a physical simulation environment for testing autonomous driving tasks under defected road conditions. Experimental results demonstrate that pre-training baseline perception networks with our synthetic data significantly improves their performance, addressing the scarcity of annotated real-world road defect data and supporting the development of robust inspection models. Additionally, decision-making strategies for navigating defected roads are evaluated. Common 2D path planning algorithms are analyzed for bypassing or gliding over road defects. When defects are unavoidable, vibration measurements under different traversal speeds reveal that optimal speed strategies vary based on the severity of the defect.
Despite these advancements, several challenges remain. The sim-to-real gap is inevitable due to information loss during the rendering process, leading to synthetic data with artificial appearances. As a result, synthetic data is primarily limited to pre-training and must be supplemented with real-world data for fine-tuning. Future work could explore the use of explicit representations, such as 3D Gaussian Splatting (3DGS) [77], to replace the current rendering process with photorealistic novel view synthesis, while preserving the explicit geometric accuracy required for physical simulation.

### v2

@@ III-B Digital Road Twin Generator
with similar solutions for and . With all intersecting faces removed, the road defect models are positioned within the holes of the planar road mesh. The planar road polygons are then re-organized, typically through triangulation, which iteratively divides edges, samples vertices, and generates faces. However, triangle mismatches may occur at the boundaries if adjacent meshes are reconstructed separately. For better realism and diversity, real-world road surface elevations are restored from the library. The height of each vertex is efficiently queried through a k-d Tree built from the surface model. The final integrated 3D mesh is then split into individual assets, as required by the simulator, to provide semantic references.
@@
IV Experimental Results
IV-A Experimantal Setups
The sensor setup (Fig. 1) is mounted on a remotely-controlled vehicle for continuous data collection over campus roads. Images and LiDAR point clouds are captured at 10Hz, while coarse GNSS coordinates are recorded to enable defect localization and support large-scale maintenance in the future. The GNSS module also outputs signals for hardware-level synchronization. The sensor poses are estimated using FAST-LIO [fastlio], which builds and updates a coarse-grained large-scale 3D road point cloud. The points belonging to drivable areas are filtered by Grounded-SAM, while stereo matching is applied to generate dense 3D reconstructions of detected defects [fan2019pothole]. Based on sensor data collected from real-world defected roads, our hierarchical road model creator generates a library containing 53 distinct road defect models. These models are integrated into scenes in the simulator, replacing the original planar road surfaces. For example, roads in CARLA Town01 are replaced by 646 synthetic road segments and 1,000 defect instances using our digital road twin generator. In the simulation, an autopiloted vehicle navigates through road defect scenes populated with randomly placed surrounding vehicles and pedestrians. To mitigate the sim-to-real gap in the captured data, we introduce controlled randomness into the data acquisition pipeline. Specifically, sensor poses are sampled with random viewing angles rather than fixed trajectories, reducing view-dependent bias. Road surfaces and defects are rendered with diverse materials, including asphalt, cobblestones, and cement. Dynamic weather conditions are continuously simulated to enhance variability. To demonstrate the effectiveness of our synthetic data, we selected models from MMSegmentation 111MMSegmentation: https://github.com/open-mmlab/mmsegmentation with diverse backbones to ensure architectural diversity. For stereo matching, we employed representative state-of-the-art general-purpose models to provide a fair and rigorous baseline.
IV-B Perception Tasks
IV-B1 Semantic Segmentation
We first evaluate single-modal semantic segmentation networks using real-world RGB images. UDTIRI [guo2024udtiri], a real-world road defect dataset containing 1,000 semantically annotated images, is used for evaluation. All networks are pre-trained for 40,000 epochs by controlled ratios of synthetic data, followed by fine-tuning on 600 images from the real-world dataset. Network configurations follow the default settings of the MMSegmentation benchmark. To ensure fair comparisons, all hyperparameters are kept identical across experiments, with the only difference being the data for pretraining. Qualitative and quantitative results are presented in Fig. 6 and Table I, respectively. It is evident that pre-training with our synthetic data significantly enhances segmentation accuracy, providing more precise boundaries and reducing the occurrence of false-positive regions. The results show that synthetic data positively impacts image-based road defect detection, increasing mIoU by up to 3%. This improvement alleviates the challenge posed by the limited availability of annotated road defect data for real-world inspection models.
@@ IV-B1 Semantic Segmentation
| Mask2Former | 77.08 | 79.79 | 80.07 |
| PSPNet | 71.42 | 71.44 | 74.29 |
| SegFormer | 75.16 | 75.96 | 76.94 |
| SegNext | 79.68 | 79.84 | 80.63 |
| Twins | 78.40 | 79.10 | 79.06 |
| UperNet | 70.97 | 72.54 | 73.89 |
@@
| RTFNet | 67.5 | 74.6 | 69.2 | 63.6 |
| FuseNet | 66.6 | 78.7 | 53.9 | 58.5 |
| MFNet | 72.0 | 80.1 | 90.6 | 72.3 |
| SNE-RoadSeg | 67.4 | 78.8 | 94.5 | 73.5 |
| RoadFormer | 90.3 | 94.8 | 96.1 | 90.6 |
@@
Single-modal prediction results are often vulnerable to varying weather and illumination. In contrast, multi-modal fusion is widely regarded as a solution to enhance perception robustness by utilizing auxiliary information sources. However, capturing these modalities remains challenging. Leveraging the G-buffer in the rendering pipeline, we are able to evaluate auxiliary modalities for road inspection, focusing on image segmentation for drivable areas, road defects, and backgrounds, as shown in Fig. 7. As shown in Table II, each additional modality is fused with RGB images in an RGB+X format. RGB+Depth provides the most generalized results, outperforming those trained using only RGB images by 4.5% to 12.1%. It also generates the most accurate true-positive predictions, demonstrating robust performance across all cases. This result is intuitive, as depth images directly capture the surface structure. Normal maps, which are highly sensitive to object surfaces, improve performance by 1.7% to 27.1% across most networks, except for FuseNet. Although useful in many cases, this sensitivity also introduces noise, leading to convergence issues. Simple fusion of event images across network channels is unstable. Although event images improve SNE-RoadSeg by 6.1%, they degrade FuseNet’s performance by 8.1%. This instability is likely due to event data being influenced by the vehicle’s driving speed, whereas RGB images assume static objects during exposure. To effectively utilize this modality, additional speed priors are required.
IV-B2 Stereo Matching
For stereo matching evaluation, we compare five SoTA networks to demonstrate the improvements enabled by our system. Similar to the segmentation experiments, all hyperparameters are kept fixed according to the released default settings, with the only difference being whether the models are pre-trained on synthetic UDT data. Unlike individual road surface images that are available from crowd-sourcing applications, stereo images of road surface defects are exceedingly rare. To address this limitation, we fine-tune the networks using the KITTI dataset [kitti], a widely used driving dataset but does not contain road defects. Experiments are conducted on both synthetic data with ground truth annotations and the real-world Stereo-Road dataset [fan2019pothole] to evaluate zero-shot performance. Several metrics are selected for evaluation, including percentage of error pixels (PEP), which represents the percentage of incorrect disparities with respect to a tolerance of pixels, and end-point error (EPE), which measures the average disparity estimation error. Additionally, mean squared error (MSE), structural similarity index (SSIM), and peak signal-to-noise ratio (PSNR) are compared between the predicted and ground truth disparity maps. The loss function, learning rate, and optimizer settings are consistent with those reported in the corresponding publications. Table III presents the quantitative results tested on synthetic data. Nearly all networks pre-trained with our synthetic data demonstrate performance improvements. EPE is improved across all networks, with reductions ranging from 0.60 to 3.18. These results confirm that our UDT system significantly enhances model performance for stereo matching tasks.
| Method | Evaluation on Synthetic UDT Data | Evaluation on Stereo-Road Dataset (Zero-shot) | |||||||
| PEP(%) | EPE(pixel) | SSIM | MSE | PSNR | SSIM* | MSE* | PSNR* | ||
| PSMNet | 53.50 | 22.00 | 2.57 | 0.82 | 152.2 | 28.3 | 0.84 | 129.0 | 28.3 |
| PSMNet+synthetic | 7.61 | 3.59 | 0.67 | 0.90 | 93.5 | 30.9 | 0.92 | 64.7 | 31.1 |
| AANet | 37.60 | 23.7 | 4.79 | 0.85 | 118.2 | 29.7 | 0.77 | 198.7 | 25.9 |
| AANet+synthetic | 9.36 | 5.30 | 0.61 | 0.91 | 93.4 | 30.8 | 0.93 | 67.7 | 30.7 |
| LacGwc | 15.00 | 8.17 | 1.35 | 0.89 | 113.4 | 30.4 | 0.92 | 75.8 | 30.3 |
| LacGwc+synthetic | 6.11 | 2.92 | 0.55 | 0.90 | 96.0 | 30.7 | 0.93 | 63.3 | 31.1 |
| IGEV | 8.14 | 4.99 | 0.83 | 0.90 | 102.9 | 30.8 | 0.93 | 64.2 | 31.1 |
| IGEV+synthetic | 3.17 | 1.59 | 0.23 | 0.91 | 92.5 | 30.9 | 0.93 | 60.7 | 31.3 |
| ViTAStereo | 7.61 | 5.08 | 1.31 | 0.90 | 124.4 | 29.6 | 0.92 | 79.6 | 29.9 |
| ViTAStereo+synthetic | 3.14 | 1.61 | 0.22 | 0.91 | 91.6 | 30.9 | 0.93 | 60.0 | 31.3 |
IV-C Decision-Making Tasks
@@ IV-C1 Path Planning
In scenarios with ample road space and minimal opposing traffic, avoiding defects by following an alternative path is generally the preferred strategy. Unlike traditional obstacle constraints, road defects primarily impact the wheels rather than the vehicle body, allowing the vehicle to either bypass or glide over the defects. As shown in Fig. 9(a), multiple defects are distributed along the road. The drivable area and defected regions are modeled as grid maps with a resolution of 0.2m, and the vehicle model assumes a width of 2.0m and a rear-to-front wheel distance of 3.0m. Four widely used planning algorithms, A*, RRT* , Lattice, and Hybrid A*, are compared. Evaluation metrics follow those from classical obstacle avoidance tasks. Path deviation quantifies the length difference between the planned path and the shortest possible path, typically the road’s midline. Path smoothness is calculated by averaging the angle changes along the trajectory, while obstacle clearance measures the mean distance between the vehicle’s wheels and the nearest road defects. Quantitative and qualitative results of the generated paths are presented in Table IV and Fig. 9(a). Notably, A* and RRT* do not account for vehicle kinodynamic constraints, resulting in relatively poor path smoothness. However, their discrete path points remain topologically reasonable. The paths generated by Hybrid A* and Lattice glide over defects and , which is justified in the given context, respectively. Lattice generates the shortest and smoothest path, with a path deviation of 0.87 and a smoothness value of 0.047. However, because its grids and curves are pre-sampled, obstacles may not always be perfectly avoided. In contrast, Hybrid A* expands nodes iteratively at each step, achieving an obstacle clearance of 3.74, which ensures better safety distances. The trade-off, however, is that heuristic terms must be manually designed to suit specific scenarios.
IV-C2 Speed Control
In scenarios where lane changes pose a high risk, the vehicle must traverse road defects directly. In such cases, longitudinal speed becomes the only controllable variable to minimize vibrations. We customize these scenarios using our UDT system, where real-world road surface models are cropped into 40-meter sections, and integrated with randomly distributed road defects. With the aid of simulators, road defects along the wheel trajectories can be precisely located, enabling informed speed control strategies. To measure vehicle vibrations caused by defected road surfaces, collision detection is detected at the level of individual triangular faces between the road and tires during physical simulation. As shown in Fig. 9(b), six-axis IMU data is recorded as the vehicle drives at a constant speed. During the process of a wheel falling into and exiting a road defect, it can be observed that the most affected parameters are vertical acceleration , rotational velocity in roll and pitch . Subsequently, the vibration degree is defined as , where is set to 0.1 in this scenario. Quantitative results for traversing three road defects at different speeds are presented in Fig. 9(c). The results indicate that, for large and severe defects such as , slowing down is the most effective strategy for driving comfort. Interestingly, for small and shallow defects, such as , increasing speed can also reduce vibrations. Although initially counterintuitive, this observation aligns with practical experience upon closer analysis.
@@ IV-C2 Speed Control
V Discussion and Future Work
Despite these advancements, several challenges remain. First, the defect taxonomy is not yet comprehensive. The current library primarily models geometric surface deformations such as potholes and depressions, whereas finer-scale surface distresses, including cracking, raveling, and bleeding, are not yet reconstructed. Achieving a more complete taxonomy will require larger-scale data acquisition and systematic defect surveying in future deployments. Second, loss of geometric and texture fidelity is inevitable during the reconstruction process, leading to domain discrepancy between reconstructed virtual environments and real-world observations. As a result, synthetic data is primarily limited to pre-training and must be supplemented with real-world images for fine-tuning. Future work could explore domain adaptation techniques and advanced rendering methods to reduce the sim-to-real gap. Another promising direction is to adopt novel view synthesis (NVS) approaches, which learn neural scene representations directly from multi-view imagery without explicit geometric reconstruction. Unlike traditional geometry pipelines, NVS has the potential to mitigate fidelity loss during reconstruction and deliver more photorealistic renderings for downstream tasks. Another limitation concerns deployment. Due to constraints of our single prototype and the campus-scale environment, we have not yet explored large-scale, in-the-wild deployment scenarios that are essential for real-world road inspection. In the future, the generated synthetic data can be exploited for model fine-tuning and knowledge distillation, thereby enabling the development of lightweight models capable of real-time operation in large-scale road monitoring applications.
VI Conclusion
This letter presents a UDT system that exploits reality–virtual interconnections of road entities for intelligent road inspection. The system captures data from defective roads using a custom sensor setup, constructs detailed defect and surface models through a hierarchical road model creator, and integrates them into simulation scenes via a digital road twin generator to recover simplified planar road assets. The proposed system provides unlimited, multi-modal, and well-annotated data for training road inspection networks, as well as a physical simulation environment for testing autonomous driving tasks under defected road conditions. Experiments show that pre-training semantic segmentation networks with our synthetic data substantially boosts performance, alleviating the shortage of annotated real-world defect data. The precise modeling of road defects enables both trajectory planning to bypass defects and speed control to traverse them smoothly, balancing safety and comfort. Finally, we discuss current limitations in addressing the sim-to-real gap and challenges in in-the-wild deployment, outlining future directions for improved realism and large-scale applications.